OpenStudy (anthony0223990):

Gold in the form of solid cylinders will be molded into solid blocks. Each cylinder has a diameter of 6.19 cm and a height of 6 cm. Each solid block will have a volume of 360 cm3. How many cylinders of gold will it take to mold one block of gold?

OpenStudy (anthony0223990):

Fan me drummer girl

OpenStudy (drummergirl3):

Volume of each cylinder = PI r² height V = PI * 3.095² * 6 cm V = 180.5603674113 Each block = 360 cm³ # of cylinders = 360 cm³ / 180.5603674113 = 360 / 180.5603674113 = 1.9938 cylinders (or rounded = ) 2 cylinders.

OpenStudy (drummergirl3):

you find the volume of each cylinder and see how many cylinders gives you 360 cm^3 of volume

OpenStudy (anthony0223990):

Your cute and smart damn

OpenStudy (drummergirl3):

hahaha thanks you i love math <4